# VisMem: Latent Vision Memory Unlocks Potential of Vision-Language Models

### v1

@@ 4.1 Settings
Benchmarks. We select 12 benchmarks to comprehensively evaluate three main abilities of VLMs, i.e., understanding, reasoning and generation [li2024survey]. These benchmarks include: (1) Understanding: MMStar [chen2024we], MMVet [yu2024mmvet], MMT [ying2024mmt], BLINK [fu2024blink], MuirBench [wang2025muirbench]; (2) Reasoning: MMMU [yue2024mmmu], LogicVista [xiao2024logicvista], Math-V [wang2024measuring], MV-Math [wang2025mv]; (3) Generation: HallBench [guan2024hallusionbench], MultiTrust [zhang2024multitrust], MMVU [liu2025unveiling]. Details are in Appendix 8.2.
Baselines. We compare our VisMem against 15 baselines, falling into four categories: (a) direct training methods: SFT, Visual-RFT [liu2025visual], VLM-R1 [shen2025vlm], Vision-R1 [huang2025vision] and PAPO [wang2025perception]; (b) image-level methods: GRIT [fan2025grit], Sketchpad [hu2024visual], MVoT [li2025imagine], OpenThinkImg [su2025openthinkimg] and DeepEyes [zheng2025deepeyes]; (c) token-level methods: Scaffold [lei2024scaffolding], MINT-CoT [chen2025mint], ICoT [gao2025interleaved], and VPT [yu2025introducing]; (d) latent space methods: Mirage [yang2025machine]. Details are in Appendix 8.3.
Implementation Details. All experiments (except for Tab. 2) are implemented on Qwen2.5-VL-7B [bai2025qwen2] based on 8 NVIDIA H200 141G GPUs. The length of memory query $K$ is set to 8, and the lengths of short-term $N_{s}$ and long-term latent vision memory $N_{l}$ are 8 and 16, respectively. More implementation details are listed in Appendix 8.4.
@@ 4.2 Main Results
Table 2: Results on nine base models with various sizes and sources, including Qwen2.5-VL-3B/7B/32B [bai2025qwen2], LLaVA-OV-1.5-4B/8B [an2025llava], InternVL-3.5-4B/8B/14B/38B [wang2025internvl3]. $\uparrow$ indicates the performance enhancement compared with the base model.
@@
Star Vet Bench Vista -V -Math Bench Trust
Qwen2.5-VL-3B [bai2025qwen2] 52.9 61.5 49.8 46.0 46.1 52.6 39.7 19.8 13.2 46.3 56.9 48.4
@@
[Enh.1] VisMem enables advanced and comprehensive visual capabilities. As presented in Tab. 1, our proposed method demonstrates distinct superiority over other baseline models. Compared with the vanilla model, VisMem achieves a notable average improvement of 11.8% across all benchmarks. When compared with the top three baselines (i.e., Vision-R1 [huang2025vision], VLM-R1 [shen2025vlm], and OpenThinkImg [su2025openthinkimg]), our method still maintains improvements of 3.1%, 4.5%, and 4.8%, respectively. Furthermore, it consistently enhances performance across the three core domains of visual tasks, namely, understanding, reasoning, and generation. Our latent vision memory mechanism yields comprehensive enhancements in visual capabilities, with specific gains of +8.9% in visual understanding, +16.4% in reasoning, and +10.6% in generation, relative to the vanilla model. It is also noteworthy that direct RL-based methods (e.g., VLM-R1 [shen2025vlm] and Vision-R1 [huang2025vision]) also achieve relatively better performance than most other paradigms. However, this approach of directly modifying parameters relies on incremental parameter updates, which may lead to the overwriting of prior general knowledge and result in catastrophic forgetting.
@@
As illustrated in Tab. 5 and 6, we conduct additional evaluations on selected subsets of MuirBench [wang2025muirbench] and LogicVista [xiao2024logicvista]. Endowed with short- and long-term vision memory, our VisMem outperforms all baseline methods by a substantial margin in tasks demanding fine-grained visual evidence, such as counting (+7.0%), visual retrieval (+9.4%), and grounding (13.1%), while also yielding notable improvements in visual reasoning tasks, including inductive (+5.7%) and deductive (+7.1%) learning.
@@
Figure 3: Results of the cross-domain generalization study. Models are only trained on Visual CoT [shao2024visual] and Mulberry [yao2024mulberry]. Dashed bar indicates the results with full training data.
@@
[Enh.2] VisMem showcases great cross-domain generalization. To evaluate the cross-domain generalization capability of our model, specifically whether its stored latent visual memory can transfer across diverse unseen tasks, we exclusively train our VisMem and comparative baseline models on two datasets: Visual CoT [shao2024visual] and Mulberry [yao2024mulberry], then subsequently assess their performance on four unseen target benchmarks. As demonstrated in Fig. 3, 7, and Tab. 7, VisMem not only consistently achieves significant performance gains on out-of-domain tasks (+6.9% on MMVet [yu2024mmvet], +9.1% on MuirBench [wang2025muirbench], +20.2% on MV-Math [wang2025mv], and +9.9% on MultiTrust [zhang2024multitrust]), but also maintains leading performance relative to all baselines. Notably, our method outperforms the second-ranked model by a substantial margin of 2.7–6.8% across all four benchmarks, while narrowing the performance gap relative to results obtained with full training data. This observation underscores its robust cross-domain knowledge transfer capability.
@@
Figure 4: Results of four-stage continual learning on MMVet [yu2024mmvet]. Stage 0 only includes itself, while stage 1, 2, 3 sequentially train models on different additional training data combinations.
@@
[Enh.3] VisMem alleviates catastrophic forgetting. As illustrated in Fig. 4, 8, and Tab. 8, we conduct sequential training of the models across four stages, with performance assessed on MMVet [yu2024mmvet] after each stage. At stage 0, the model was trained exclusively on the base task, and in subsequent stages, we incrementally incorporated selected benchmarks into the training process. From the continual learning results, our VisMem demonstrates significantly stronger knowledge retention capabilities. Although direct training paradigms yield relatively excellent overall performance in offline learning tasks with once-off training, they suffer from severe catastrophic forgetting. For instance, SFT exhibits over 10% performance degradation throughout the training process, the highest among all baselines. Additionally, at stage 0, VLM-R1 [shen2025vlm] and Vision-R1 [liu2025visual] achieve performance improvements of 11.8% and 10.9% respectively compared to the vanilla model, however, these improvements are retained by less than 0.5% at stage 4. In contrast, our method effectively mitigates catastrophic forgetting, exhibiting the smallest performance gap relative to original full-data training among all baselines. It is further worth noting that our latent vision memory enhances performance at stages 1 and 3 without any degradation, reflecting superior cross-task generalization.
@@ 4.3 Additional Analyses
[Obs.1] VisMem is robustly compatible across various base models. As detailed in Tab. 2 and Fig. 11, to evaluate the generalizability of our approach across diverse base models, we assess nine widely used base models, encompassing Qwen2.5-VL-3B/32B [bai2025qwen2], LLaVA-OV-1.5-4B/8B [an2025llava], InternVL-3.5-4B/8B/14B/38B [wang2025internvl3], with parameter scales ranging from 3B to 38B. The results indicate that our latent vision memory paradigm exhibits strong compatibility across various models, yielding significant performance improvements across most visual tasks.
[Obs.2] The memory invocations are dynamic and self-adaptive. To elaborate on the effectiveness of our dual latent memory system, we characterize the properties of the short- and long-term memories it forms. As illustrated in Fig. 5, we first analyze the type-specific invocation ratios and their relative positions within the output sequence across four benchmarks. In summary, invocation ratios are self-adaptive across tasks, while both memory types exhibit a dynamic downward trend in invocation frequency throughout the output sequence. Task-specific comparisons in Fig. 9 further reveal that short-term latent memories are invoked more frequently to retrieve fine-grained details during visual information acquisition and understanding, particularly in multi-image scenarios, such as MuirBench [wang2025muirbench]. Conversely, long-term latent vision memories play a more critical role in reasoning, e.g., in MV-Math [wang2025mv], by providing abstract semantic knowledge relevant to the current task. Furthermore, Tab. 5 and 6, which detail the sub-task performance of MuirBench [wang2025muirbench] and LogicVista [xiao2024logicvista] respectively, further illustrate that short-term and long-term latent visual memories are complementary. Their dynamic invocation yields superior performance compared to relying on a single memory type or the absence of vision memory.
(a) MMVet [yu2024mmvet]
(b) MuirBench [wang2025muirbench]
(c) MV-Math [wang2025mv]
(d) MultiTrust [zhang2024multitrust]
@@
(a) MMVet [yu2024mmvet]
(b) MuirBench [wang2025muirbench]
(c) MV-Math [wang2025mv]
(d) MultiTrust [zhang2024multitrust]
@@ 6 Theoretical Foundations
As the mainstream position in anthropological cognitive psychology since the 20th century, short-term memory and long-term memory are two distinct storage systems that can be differentiated based on their functional and neural underpinnings [baddeley2012working, norris2017short]. Specifically, the Dennis Norris Theory [norris2017short] proposes that short-term memory requires processing new visual information, temporarily storing multiple tokens, and enabling variable signals. It relies neurologically on vision-specific brain regions, e.g., the visual cortex and the posterior superior temporal lobe associated with verbal short-term memory), exhibiting visual dominance; long-term memory, however, centers on abstract semantic representations and relies on semantic-related brain regions like the medial temporal lobe and mid-temporal lobe.
Thus, we propose a framework termed VisMem to invoke dual short and long latent memory during the token-by-token autoregressive generation. Aligned with Dennis Norris Theory [norris2017short], we instantiate these roles in a VLM backbone via latent vision memory invocation and latent vision memory formation, which together produce distinct short and long latent memory tokens and integrate them into the generation stream of the model.
@@ 7.2 Training Recipe
As mentioned in Sec. 3.4, we design a two-stage training pipeline: at the first stage, the main objective is to optimize the memory formation process (see Eq. 7); at the second stage, the main objective is to optimize the memory invocation (see Eq. 8). We update the models based on reinforcement learning, i.e., GRPO strategy [shao2024deepseekmath]. Specifically, for each instruction-vision pair $\left(I,V\right)$, the policy model $\mathcal{P}$ generates a group of $G$ distinct candidate trajectories, termed as $\mathcal{T}=\left\{\tau_{1},\dots,\tau_{G}\right\}$. For each trajectory, we utilize a $S\left(\cdot\right)$ to quantify the performance. Then, a group-relative baseline is calculated via averaging and standardizing all trajectories within the candidate group $G$:
@@ 8.1 Training Data
During the two-stage training procedure, we use the same training data to optimize both the memory invocation and memory formation in the latent vision memory system. Initially, we include the training split dataset of the selected benchmarks and retain their original data division. For benchmarks without a training phase, we use them solely for evaluation. Additionally, we incorporate the Visual CoT [shao2024visual] and Mullberry [yao2024mulberry], improving the reasoning abilities.
@@ 8.2 Benchmarks
  • MMStar [chen2024we] is a high-quality vision-centric benchmark meticulously curated by human experts. This benchmark assesses 6 core capabilities across 18 detailed axes of visual understanding.
  • MMVet [yu2024mmvet] establishes 6 core visual understanding capabilities and investigates 16 critical integrations derived from their combinations. It uses an evaluator tailored for open-ended outputs.
  • MMT [ying2024mmt] consists of carefully curated multi-choice visual questions, covering 32 core meta-tasks and 162 subtasks within the field of visual understanding.
  • BLINK [fu2024blink] reconstructs 14 classic computer vision tasks into multiple-choice questions. Each question is paired with either single or multiple images and supplemented with visual prompting.
  • MuirBench [wang2025muirbench] covers 12 diverse multi-image tasks, which involve 10 categories of multi-image relations. Each standard instance is paired with an unanswerable variant that differs only minimally in semantics.
  • MMMU [yue2024mmmu] comprises meticulously curated visual questions sourced from college exams, quizzes, and textbooks spanning 30 subjects and 183 subfields, which focus on advanced reasoning grounded in domain-specific knowledge.
  • LogicVista [xiao2024logicvista] evaluates general logical cognition abilities across 5 logical reasoning tasks, which encompass 9 distinct capabilities. Each question is annotated with the correct answer and the human-written reasoning behind the selection.
  • Math-V [wang2024measuring] is comprised of high-quality mathematical problems with visual contexts, drawn from real mathematical competitions. It spans 16 distinct mathematical disciplines and is categorized into 5 difficulty levels, evaluating the mathematical reasoning.
  • MV-Math [wang2025mv] is a dataset comprising mathematical problems, integrating multiple images interleaved with text, and detailed annotations. It features multiple-choice, free-form, and multi-step questions across 11 subject areas at 3 difficulty levels.
  • HallBench [guan2024hallusionbench] consists of images paired with questions, designed by human experts to assess the hallucination level of generation.
  • MultiTrust [zhang2024multitrust] covers five primary aspects: truthfulness, safety, robustness, fairness, and privacy, evaluating the trustworthiness of generation.
  • MMVU [liu2025unveiling] encompasses 12 categories, and designs evaluation metrics that measure the quality and error degree of generation.
@@
optimizeroptimizer AdamW [loshchilov2017decoupled]
@@
Table 5: Results on 9 selected subsets of MuirBench [wang2025muirbench]. We compare our VisMem with the second and third best scored counterparts, and separately use the short or long latent memory to assess the improvements of each.
@@
Vanilla [bai2025qwen2] 44.1 34.2 80.9 70.5 53.2 52.9 82.4 53.7 76.1
VLM-R1 [shen2025vlm] 52.5 38.1 83.6 73.5 58.1 55.1 86.8 56.7 79.4
Vision-R1 [huang2025vision] 53.8 39.2 84.5 73.1 57.4 57.2 87.4 57.9 78.9
@@
Table 6: Results on 10 selected subsets (5 reasoning skills and 5 capabilities) of LogicVista [xiao2024logicvista]. We compare our VisMem with the second and third best scored counterparts, and separately use the short or long latent memory to assess the improvements of each.
@@
Vanilla [bai2025qwen2] 44.6 45.0 39.7 37.9 48.7 30.1 32.5 41.6 34.4 36.8
VLM-R1 [shen2025vlm] 53.7 52.7 45.8 44.1 57.3 35.8 42.8 49.0 46.5 52.6
Vision-R1 [huang2025vision] 53.5 51.4 46.7 44.8 58.9 36.5 43.6 49.7 48.2 53.8

### v2

@@ 4.1 Settings
Benchmarks. We select 12 benchmarks to comprehensively evaluate three main abilities of VLMs, i.e., understanding, reasoning and generation [31]. These benchmarks include: (1) understanding: MMStar [7], MMVet [76], MMT [73], BLINK [15], MuirBench [57]; (2) reasoning: MMMU [79], LogicVista [67], MathVista [37], MV-Math [60]; (3) generation: HallBench [19], MultiTrust [82], MMVU [34]. Details are in Appendix 8.2.
Baselines. We compare our VisMem against 15 baselines, falling into four categories: (a) direct training methods: SFT, Visual-RFT [35], VLM-R1 [44], Vision-R1 [26] and PAPO [66]; (b) image-level methods: GRIT [13], Sketchpad [24], MVoT [29], OpenThinkImg [49] and DeepEyes [87]; (c) token-level methods: Scaffold [28], MINT-CoT [9], ICoT [16], and VPT [75]; (d) latent space methods: Mirage [70]. Details are in Appendix 8.3.
Implementation Details. All experiments (except for Tab. 2) are implemented on Qwen2.5-VL-7B [4] based on 8 NVIDIA H200 141G GPUs. The length of memory query $K$ is set to 8, and the lengths of short-term $N_{s}$ and long-term latent vision memory $N_{l}$ are 8 and 16, respectively. More implementation details are listed in Appendix 8.4.
@@ 4.2 Main Results
Table 2: Results on nine base models with various sizes and sources, including Qwen2.5-VL-3B/7B/32B [4], LLaVA-OV-1.5-4B/8B [1], InternVL-3.5-4B/8B/14B/38B [63]. $\uparrow$ indicates the performance enhancement compared with the base model.
@@
Star Vet Bench Vista Vista -Math Bench Trust
Qwen2.5-VL-3B [4] 52.9 61.5 49.8 46.0 46.1 52.6 39.7 61.0 13.2 46.3 56.9 48.4
@@
[Enh.1] VisMem enables advanced and comprehensive visual capabilities. As presented in Tab. 1, our proposed method demonstrates distinct superiority over other baseline models. Compared with the vanilla model, VisMem achieves a notable average improvement of 11.0% across all benchmarks. When compared with the top three baselines (i.e., Vision-R1 [26], VLM-R1 [44], and OpenThinkImg [49]), our method still maintains improvements of 3.0%, 4.2%, and 4.9%, respectively. Furthermore, it consistently enhances performance across the three core domains of visual tasks, namely, understanding, reasoning, and generation. Our latent vision memory mechanism yields comprehensive enhancements in visual capabilities, with specific gains of +8.9% in visual understanding, +14.4% in reasoning, and +10.6% in generation, relative to the vanilla model. It is also noteworthy that direct RL-based methods (e.g., VLM-R1 [44] and Vision-R1 [26]) also achieve relatively better performance than most other paradigms. However, this approach of directly modifying parameters relies on incremental parameter updates, which may lead to the overwriting of prior general knowledge and result in catastrophic forgetting.
@@
As illustrated in Tab. 5 and 6, we conduct additional evaluations on selected subsets of MuirBench [57] and LogicVista [67]. Endowed with short- and long-term vision memory, our VisMem outperforms all baseline methods by a substantial margin in tasks demanding fine-grained visual evidence, such as counting (+7.0%), visual retrieval (+9.4%), and grounding (13.1%), while also yielding notable improvements in visual reasoning tasks, including inductive (+5.7%) and deductive (+7.1%) learning.
@@
Figure 3: Results of the cross-domain generalization study. Models are only trained on Visual CoT [42] and Mulberry [71]. Dashed bar indicates the results with full training data.
@@
[Enh.2] VisMem showcases great cross-domain generalization. To evaluate the cross-domain generalization capability of our model, specifically whether its stored latent visual memory can transfer across diverse unseen tasks, we exclusively train our VisMem and comparative baseline models on two datasets: Visual CoT [42] and Mulberry [71], then subsequently assess their performance on four unseen target benchmarks. As demonstrated in Fig. 3, 7, and Tab. 7, VisMem not only consistently achieves significant performance gains on out-of-domain tasks (+6.9% on MMVet [76], +9.1% on MuirBench [57], +20.2% on MV-Math [60], and +9.9% on MultiTrust [82]), but also maintains leading performance relative to all baselines. Notably, our method outperforms the second-ranked model by a substantial margin of 2.7–6.8% across all four benchmarks, while narrowing the performance gap relative to results obtained with full training data. This observation underscores its robust cross-domain knowledge transfer capability.
@@
Figure 4: Results of four-stage continual learning on MMVet [76]. Stage 0 only includes itself, while stage 1, 2, 3 sequentially train models on different additional training data combinations.
@@
[Enh.3] VisMem alleviates catastrophic forgetting. As illustrated in Fig. 4, 8, and Tab. 8, we conduct sequential training of the models across four stages, with performance assessed on MMVet [76] after each stage. At stage 0, the model was trained exclusively on the base task, and in subsequent stages, we incrementally incorporated selected benchmarks into the training process. From the continual learning results, our VisMem demonstrates significantly stronger knowledge retention capabilities. Although direct training paradigms yield relatively excellent overall performance in offline learning tasks with once-off training, they suffer from severe catastrophic forgetting. For instance, SFT exhibits over 10% performance degradation throughout the training process, the highest among all baselines. Additionally, at stage 0, VLM-R1 [44] and Vision-R1 [35] achieve performance improvements of 11.8% and 10.9% respectively compared to the vanilla model, however, these improvements are retained by less than 0.5% at stage 4. In contrast, our method effectively mitigates catastrophic forgetting, exhibiting the smallest performance gap relative to original full-data training among all baselines. It is further worth noting that our latent vision memory enhances performance at stages 1 and 3 without any degradation, reflecting superior cross-task generalization.
@@ 4.3 Additional Analyses
[Obs.1] VisMem is robustly compatible across various base models. As detailed in Tab. 2 and Fig. 11, to evaluate the generalizability of our approach across diverse base models, we assess nine widely used base models, encompassing Qwen2.5-VL-3B/32B [4], LLaVA-OV-1.5-4B/8B [1], InternVL-3.5-4B/8B/14B/38B [63], with parameter scales ranging from 3B to 38B. The results indicate that our latent vision memory paradigm exhibits strong compatibility across various models, yielding significant performance improvements across most visual tasks.
[Obs.2] The memory invocations are dynamic and self-adaptive. To elaborate on the effectiveness of our dual latent memory system, we characterize the properties of the short- and long-term memories it forms. As illustrated in Fig. 5, we first analyze the type-specific invocation ratios and their relative positions within the output sequence across four benchmarks. In summary, invocation ratios are self-adaptive across tasks, while both memory types exhibit a dynamic downward trend in invocation frequency throughout the output sequence. Task-specific comparisons in Fig. 9 further reveal that short-term latent memories are invoked more frequently to retrieve fine-grained details during visual information acquisition and understanding, particularly in multi-image scenarios, such as MuirBench [57]. Conversely, long-term latent vision memories play a more critical role in reasoning, e.g., in MV-Math [60], by providing abstract semantic knowledge relevant to the current task. Furthermore, Tab. 5 and 6, which detail the sub-task performance of MuirBench [57] and LogicVista [67] respectively, further illustrate that short-term and long-term latent visual memories are complementary. Their dynamic invocation yields superior performance compared to relying on a single memory type or the absence of vision memory.
(a) MMVet [76]
(b) MuirBench [57]
(c) MV-Math [60]
(d) MultiTrust [82]
@@
(a) MMVet [76]
(b) MuirBench [57]
(c) MV-Math [60]
(d) MultiTrust [82]
@@ 6 Theoretical Foundations
As the mainstream position in anthropological cognitive psychology since the 20th century, short-term memory and long-term memory are two distinct storage systems that can be differentiated based on their functional and neural underpinnings [3, 38]. Specifically, the Dennis Norris Theory [38] proposes that short-term memory requires processing new visual information, temporarily storing multiple tokens, and enabling variable signals. It relies neurologically on vision-specific brain regions, e.g., the visual cortex and the posterior superior temporal lobe associated with verbal short-term memory), exhibiting visual dominance; long-term memory, however, centers on abstract semantic representations and relies on semantic-related brain regions like the medial temporal lobe and mid-temporal lobe.
Thus, we propose a framework termed VisMem to invoke dual short and long latent memory during the token-by-token autoregressive generation. Aligned with Dennis Norris Theory [38], we instantiate these roles in a VLM backbone via latent vision memory invocation and latent vision memory formation, which together produce distinct short and long latent memory tokens and integrate them into the generation stream of the model.
@@ 7.2 Training Recipe
As mentioned in Sec. 3.4, we design a two-stage training pipeline: at the first stage, the main objective is to optimize the memory formation process (see Eq. 7); at the second stage, the main objective is to optimize the memory invocation (see Eq. 8). We update the models based on reinforcement learning, i.e., GRPO strategy [43]. Specifically, for each instruction-vision pair $\left(I,V\right)$, the policy model $\mathcal{P}$ generates a group of $G$ distinct candidate trajectories, termed as $\mathcal{T}=\left\{\tau_{1},\dots,\tau_{G}\right\}$. For each trajectory, we utilize a $S\left(\cdot\right)$ to quantify the performance. Then, a group-relative baseline is calculated via averaging and standardizing all trajectories within the candidate group $G$:
@@ 8.1 Training Data
During the two-stage training procedure, we use the same training data to optimize both the memory invocation and memory formation in the latent vision memory system. Initially, we include the training split dataset of the selected benchmarks and retain their original data division. For benchmarks without a training phase, we use them solely for evaluation. Additionally, we incorporate the Visual CoT [42] and Mullberry [71], improving the reasoning abilities.
@@ 8.2 Benchmarks
  • MMStar [7] is a high-quality vision-centric benchmark meticulously curated by human experts. This benchmark assesses 6 core capabilities across 18 detailed axes of visual understanding.
  • MMVet [76] establishes 6 core visual understanding capabilities and investigates 16 critical integrations derived from their combinations. It uses an evaluator tailored for open-ended outputs.
  • MMT [73] consists of carefully curated multi-choice visual questions, covering 32 core meta-tasks and 162 subtasks within the field of visual understanding.
  • BLINK [15] reconstructs 14 classic computer vision tasks into multiple-choice questions. Each question is paired with either single or multiple images and supplemented with visual prompting.
  • MuirBench [57] covers 12 diverse multi-image tasks, which involve 10 categories of multi-image relations. Each standard instance is paired with an unanswerable variant that differs only minimally in semantics.
  • MMMU [79] comprises meticulously curated visual questions sourced from college exams, quizzes, and textbooks spanning 30 subjects and 183 subfields, which focus on advanced reasoning grounded in domain-specific knowledge.
  • LogicVista [67] evaluates general logical cognition abilities across 5 logical reasoning tasks, which encompass 9 distinct capabilities. Each question is annotated with the correct answer and the human-written reasoning behind the selection.
  • MathVista [59] unifies the challenges of heterogeneous mathematical and visual tasks, which are curated from math-oriented multimodal datasets.
  • MV-Math [60] is a dataset comprising mathematical problems, integrating multiple images interleaved with text, and detailed annotations. It features multiple-choice, free-form, and multi-step questions across 11 subject areas at 3 difficulty levels.
  • HallBench [19] consists of images paired with questions, designed by human experts to assess the hallucination level of generation.
  • MultiTrust [82] covers five primary aspects: truthfulness, safety, robustness, fairness, and privacy, evaluating the trustworthiness of generation.
  • MMVU [34] encompasses 12 categories, and designs evaluation metrics that measure the quality and error degree of generation.
@@
optimizeroptimizer AdamW [36]
@@
Table 5: Results on 9 selected subsets of MuirBench [57]. We compare our VisMem with the second and third best scored counterparts, and separately use the short or long latent memory to assess the improvements of each.
@@
Vanilla [4] 44.1 34.2 80.9 70.5 53.2 52.9 82.4 53.7 76.1
VLM-R1 [44] 52.5 38.1 83.6 73.5 58.1 55.1 86.8 56.7 79.4
Vision-R1 [26] 53.8 39.2 84.5 73.1 57.4 57.2 87.4 57.9 78.9
@@
Table 6: Results on 10 selected subsets (5 reasoning skills and 5 capabilities) of LogicVista [67]. We compare our VisMem with the second and third best scored counterparts, and separately use the short or long latent memory to assess the improvements of each.
@@
Vanilla [4] 44.6 45.0 39.7 37.9 48.7 30.1 32.5 41.6 34.4 36.8
VLM-R1 [44] 53.7 52.7 45.8 44.1 57.3 35.8 42.8 49.0 46.5 52.6
Vision-R1 [26] 53.5 51.4 46.7 44.8 58.9 36.5 43.6 49.7 48.2 53.8